# Planet formation regulated by galactic-scale interstellar turbulence

### v2

@@ 2.2 Properties of local star forming regions
Motivated by this result, hereafter we we will truncate the GMC mass function above $500\,M_{\odot}$. We will also exclude regions with masses $<10M_{\odot}$, for which we expect the turbulent velocity to approach the sound speed ($\mathcal{M}\sim 1$). In principle, they may still be accurately sampled by the excursion set formalism if the density distribution does not diverge greatly from log-normal. However, we cannot choose an arbitrary small lower mass limit, since for much lower mass regions there is insufficient mass to form stellar mass objects. Experimentation shows that our results are not strongly dependent on our choice of lower mass limit.
@@ 2.3 Collapse of the giant molecular cloud
In the early stage after a star forms, it occupies the local star forming region that is undergoing gravitational collapse. Since we are interested in the local density and velocity evolution throughout a star’s lifetime, we require a simple model to follow the collapse of the GMC and formation of stars. Once the GMC becomes gravitationally unstable, it collapses on a free-fall time-scale. It does not do so monolithically, but because smaller sub-regions of the cloud may be higher density (and therefore have shorter free-fall time-scales), they fragment first and continue to accrete from their parent structures (Hoyle, 1953). Simultaneously, the cloud is exposed to heating by feedback processes from the stars as they form. A wide range of analytic and semi-analytic approaches have been developed to describe this process (Vázquez-Semadeni et al., 2019). We are helped in this context in our interest in low mass star forming regions that are locally well sampled. Since we do not expect many massive stars or supernovae among these local regions, we will hereafter ignore feedback processes and exclusively consider gravitational collapse.
We adopt the approach of Girichidis et al. (2014) in modelling the rate of star formation during the collapse of the GMC. The approach of those authors is attractive for its simplicity, while also retaining a self-consistent treatment of the gas density distribution as it evolves under gravity. We briefly review the approach here, referring the reader to the original work for details.
We consider a spherical region undergoing free-fall collapse from rest, which may be contained within a larger gravitationally unstable cloud. Girichidis et al. (2014) write an invertible, analytic approximation for the density of such a region as a function of time. Given an analytically integrable initial density distribution, there is a corresponding analytic expression for the fraction of an initial gas reservoir that collapses beyond any given threshold as a function of time. The initial density distribution in our collapsing region is log-normal, with a dispersion determined by the Mach number on the scale of collapse. We choose the critical density to be such that the local free-fall time is shorter than $0.1$ Myr. Given the accelerating nature of gravitational collapse, the exact threshold time is irrelevant so long as it is substantially shorter than the time-scale for the global collapse of the star forming region. The star formation rate is defined to be the rate at which material crosses this threshold density. We assume the cloud is instantly dispersed on cessation of the star formation phase, which continues until a star formation efficiency $\epsilon=0.5$ is reached, this value being commensurate with simulations of stellar core formation (Matzner & McKee, 2000). We expect stellar feedback, which we do not consider and may shut off star formation in massive clouds, to have a negligible influence on the low mass star forming regions that are our focus. High star formation efficiencies similar to those we assume are also reached for more sophisticated models of star formation in a collapsing cloud for low mass clouds (Vázquez-Semadeni et al., 2019).
@@ 2.4 Stellar dynamics in the bound cloud
Having defined a local star formation rate, we can draw formation times for an initial stellar population. Since these stars are initially bound to the cloud from which they form, we integrate the orbit of the star within the gravitational potential of the cloud, ignoring interactions with other stars. For convenience we will consider a Plummer potential with a core radius $a=R_{\mathrm{SFR}}/2.6$, such that the outer radius of the star forming region is twice the half-mass radius. Stars are drawn from within the core radius, and we draw an isotropic velocity distribution with a dispersion assuming instantaneous virial equilibrium with the cloud. To track local density and relative velocity of the gas, we assume a static, smooth Plummer density profile (as for Throop & Bally, 2008). We then evolve the star under the potential, varying the core radius as the spherical star forming region contracts until the final dispersal of the cloud (when $M_{*}=\epsilon M_{\mathrm{tot}}$).
2.5 Evolution of the turbulent medium
Throughout the evolution of the cloud, we continue to track the local density and velocity evolution of the background ISM. We do this once again following the procedure of Hopkins (2012). We assume the turbulence is globally steady-state, such that the turbulent velocity cascade is maintained outside of the collapsing regions. The probability density function for any Gaussian random variable $\delta$ under these conditions obeys a generalized Fokker-Planck equation. We can therefore redefine our trajectory at $t+\Delta t$ by computing a random walk step:
$\Delta\delta_{j}(t+\Delta t)=\Delta\delta_{j}(t)\,\exp{(-\Delta t/\tau_{\rm{t}})}+\mathcal{R}\,\sqrt{\Delta S\,(1-\exp{(-2\,\Delta t/\tau_{\rm{t}})})},$ (3)
where $\mathcal{R}$ is a Gaussian random number with unity variance. The new trajectory is the sum of the components given by equation 3. We define the time-step to be $2$ percent of the shortest turbulent time-scale in our grid. We evolve the ISM throughout the bound phase of star formation, but only use it to compute accretion rates after the dispersal of the cloud.
@@ 2.5 Evolution of the turbulent medium
Equation 3 applies both to the logarithmic density perturbations and to each component of the interstellar gas velocity. The relative velocity of the star as it moves through the unbound density field after the dispersal of the cloud is then the gas velocity at the relevant scale, minus the velocity of the star with respect to the bound cloud at the time of dispersal and the velocity of the cloud itself. The velocity of the cloud is fixed as the gas velocity at the time and scale of cloud collapse.
@@ 2.6 Accretion rate
BHL accretion has been the topic of many works (for a review, see Edgar, 2004). In the purest case, it is the gravitational focusing of the test particles (representing the ISM), resulting in a stagnation point in the wake of the star within which material is accreted. The instantaneous BHL radius within which material is captured is then defined:
@@
Streamlines with an impact parameter at the BHL radius $R_{\mathrm{BHL}}$ hit a stagnation point along the axis coincident with the relative upstream gas velocity vector. In the ideal case, in an homogeneous medium, they then accrete directly onto the star. However, in some circumstances the radius within which capture can occur can be smaller than the BHL radius. Firstly, clearly the accretion radius cannot exceed the turbulent size scale under consideration, $R_{\mathrm{t}}$. In addition, if the tidal radius $R_{\mathrm{T}}<R_{\mathrm{BHL}}$, then this radius can also limit the material that can be captured by the star. If a star is bound to a mass $M_{\mathrm{b}}$ at a radius $r_{\mathrm{b}}$, then:
Initially, this radius is determined by the collapsing cloud, where $r_{\mathrm{b}}$ is the instantaneous location of the star with respect to the centre of the region. Once the cloud has dispersed, we take the galactic values $r_{\mathrm{b}}=8$ kpc and $M_{\mathrm{b}}=10^{12}\,M_{\odot}$. Given these considerations, the accretion radius can be defined:
Material within the cross-section carved out by this radius is captured by the star. At any turbulent scale, we can estimate the instantaneous rate at which a star captures material in such a scenario (Bondi, 1952; Shima et al., 1985):
@@
It is now necessary to define the mass of a star to calculate accretion rates. Hereafter, stellar masses are always drawn from a log-uniform distribution between $0.01-3\,M_{\odot}$. This distribution is not important because we will resample when comparing to observations, except that we need to fully sample the stellar mass range relevant for drawing comparisons to protoplanetary disc surveys. Although the stellar masses do not evolve self-consistently as the stars accrete in our model, the total mass that they would gain is nearly always a small fraction of the initial stellar mass. We then draw 500 accretion histories from a representative sample of synthetic star forming regions, with a draw probability weighted by the stellar mass of the region.
@@
The resultant BHL accretion rates, normalised for the a solar mass star, are shown in Figure 2, for a subset of six stars in our sample. The phase in which the star remains within the bound gaseous reservoir typically persists for $\sim 1$ Myr. During this phase, local low velocity and high density gas results in an early phase of enhanced BHL accretion at a median rate $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-8}\,M_{\odot}$ yr-1 (Figure 2, bottom). This median rate then quickly drops to $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-10}{-}10^{-11}\,M_{\odot}$ yr-1 in the age range $1{-}3$ Myr for a solar mass star. This is considerably lower than the typical rate of accretion onto the central star, $\dot{M}_{\mathrm{acc}}\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 (Manara et al., 2023). One might therefore assume that BHL accretion no longer plays an important role for disc evolution. However, in a turbulent medium, vorticity of the in-falling material means that accretion is mediated via a (protoplanetary) disc (Krumholz et al., 2006; Kuffmeier et al., 2020). In this case, the change in accretion rate can be averaged over the time-scale $\tau_{\mathrm{acc}}$ at which angular momentum is transported within the disc. Empirically, this time-scale can be a considerable fraction of the disc lifetime (Almendros-Abad et al., 2024). In this case the median underestimates the true contribution of BHL accretion rate onto the disc to the accretion rate onto the star. By contrast, the mean accretion rate when averaged over all stars in our model is $\langle\dot{M}_{\mathrm{BHL}}\rangle\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 at $1-3$ Myr, as shown as the dashed red line in Figure 2. However, the mean overestimates the typical BHL accretion rate contribution for individual discs, since it is dominated by a small number of rapid accretors. Interpreting the contribution from BHL accretion to the stellar accretion rate for typical star-disc systems is therefore not a simple task. Compounding this problem further, discs that survive to comprise the observed sample may also preferentially be those which have been replenished by the ISM. Interpreting the role of BHL accretion for disc evolution therefore requires further calculation.
@@ 3.1 Simple disc model
To evolve the disc mass, in our fiducial model we fix the initial $M_{\mathrm{disc}}(t=0)=0$. This is equivalent to the assumption that the initial disc that forms due to core collapse is low mass, or rapidly depleted. This runs counter to the conventional assumption that the protoplanetary disc represents an evolved state of the primordial protostellar disc, which formed during the stellar core collapse. We make this decision as a ‘clean’ test of our model; we wish to explore the properties of discs arising from BHL accretion alone. Such a model requires no additional free parameters. We emphasise that no specific observational constraints contradict a low initial disc mass. Any such evidence would need to connect the primordial circumstellar material, which forms over $\sim 10^{2}$ years, to the protoplanetary disc, which in the context of this work forms over $\sim 10^{5}$ years and evolves over $\sim 10^{6}$ years. Although (massive) discs around Class 0 stars have been inferred (Maury et al., 2019; Tobin et al., 2020) interpreting these observations in terms of the exact age of the (proto-)star (down to $\sim 10^{4}$ years), the nature of the disc and the time-scale over which any such disc should survive is not simple. Since few, if any, unambiguous observations exist constraining the properties of primordial discs that are categorically around stars of age $\ll 10^{5}$ years, there is no reason to assume that such a primordial disc should be both long-lived and comparable in mass to the subsequently accreted one. Indeed, it would be a remarkable coincidence if BHL accretion and protostar formation contribute an equal fraction of mass to the disc population. Nonetheless, we also explore the impact of this assumption on our results below.
@@
For $\tau_{\mathrm{acc}}=M_{\mathrm{disc}}/\dot{M}_{\mathrm{acc}}$ we assume a log-uniform distribution ±$\pm 1$ dex around $\log\tau_{\mathrm{acc}}/\mathrm{1\,Myr}=-0.5$, based on visual inspection of the observed distribution (Almendros-Abad et al., 2024). Empirically, the time-scale $\tau_{\mathrm{acc}}$ is not strongly correlated with stellar mass, but may be a complex function of the disc properties or stellar age. For example, visually interpreting Figure 12 of Almendros-Abad et al. (2024), the median accretion time-scale appears to be $\tau_{\mathrm{acc,1/2}}\sim 0.1$ Myr for the young region $\rho$ Ophiuchus, while $\tau_{\mathrm{acc,1/2}}\sim 1$ Myr for the older Lupus region. Indeed, we expect rapid BHL accretion to instigate turbulence in the disc, temporarily decreasing $\tau_{\mathrm{acc}}$. However, empirically it is not clear the degree to which any variation of $\tau_{\mathrm{acc}}$ is a survival bias, discs with shorter $\tau_{\mathrm{acc}}$ having been depleted (we will show a similar evolution in our model which has fixed $\tau_{\mathrm{acc}}$ for each star below). While exploring time- or accretion-dependent $\tau_{\mathrm{acc}}$ is an interesting avenue for future exploration, we initially maintain our simplified model, subsequently exploring the potential role of the ISM in driving a time-dependent stellar accretion rate.
@@
Given the initial condition and accretion time-scale, we evolve each disc mass according to equation 9 over the entire time period under consideration. However, we assume a simple criterion for disc dispersal: $M_{\mathrm{disc}}<3\times 10^{-5}\,M_{\odot}$. We adopt this criterion when showing disc properties in subsequent analysis. This threshold is motivated empirically by the fact that few discs have masses below this value; our results are not strongly sensitive to this choice within a sensible range.

### v3

@@ 2.2.2 Mass function
Motivated by this result, hereafter we we will truncate the GMC mass function above $500\,M_{\odot}$. We will also exclude regions with masses $<10M_{\odot}$, for which we expect the turbulent velocity to approach the sound speed ($\mathcal{M}\sim 1$). In principle, they may still be accurately sampled by the excursion set formalism if the density distribution does not diverge greatly from log-normal. However, we cannot choose an arbitrary small lower mass limit, since for much lower mass regions there is insufficient mass to form stellar mass objects. Experimentation shows that our results are not strongly dependent on our choice of lower mass limit (see Section 3.4.4).
2.2.3 Caveats for the mass function
In the averaging for equation 2 we have ignored that similar mass unstable regions may be spatially correlated, since the probability of collapse on a given spatial scale is dependent on the density contribution from larger scales. However, Hopkins (2012) find that this bias is negligible for the low mass star forming regions, with $R_{\mathrm{SFR}}\ll h$. The shape of the mass function for the local, low mass GMCs relevant in the context of this work should therefore be correct. Clearly, star forming regions should still preferentially be found where the gas density is highest, while we expect using the average density is sufficient for $R_{\mathrm{obs}}\gtrsim h$.
We also highlight that our approach to computing the mass function of GMCs is not exact because we do not follow the time evolution of the interstellar medium – i.e. we follow Section 3 of Hopkins (2012) rather than Section 7, which tracks turbulent fluctuations and assembly of merger trees. In principle, in order to compute the mass function of observed star forming regions we should take into account both the variations in the turbulent timescale at different spatial scales, and how the cloud evolves after it reaches critical density. To some degree, this can be folded in by evolving the turbulent medium via perturbations to each trajectory, as described in Appendix A.3, and following Section 7 of Hopkins 2012. However, strictly this would also require self-consistently evolving the cloud, including collapse and growth. While this represents an aim for the future, we here wish to keep our model as simple as possible. We do not expect our simplification in assuming a single turbulent time-scale to strongly influence our results, primarily because the turbulent time-scale is a shallow function of $R$ ($\tau_{\mathrm{t}}\propto R^{1/2}$ for $R\ll h$), and the mass of the star forming region is a steep function of $R$ ($M\propto R^{3}$ for $R\ll h$), so that the turbulent time-scale is a very weak function of the mass of the star forming region ($\tau_{\mathrm{t}}\propto M^{1/6}$ for $R\ll h$). Indeed, we are vindicated in our assumptions in that we approximately recover $\langle\mathrm{d}N/\mathrm{d}M\rangle\propto M^{-2}$, as is found observationally across a wide range of galactic environments (e.g. Mok et al., 2020).
@@ 2.3 Collapse of the giant molecular cloud
2.3.1 Approach summary
In the early stage after a star forms, it occupies the local star forming region that is undergoing gravitational collapse. Since we are interested in the local density and velocity evolution throughout a star’s lifetime, we require a simple model to follow the collapse of the GMC and formation of stars. To do so, we first draw a representative sample of GMCs following Section 2.2, each with an associated trajectory that we will subsequently use. For each of these regions, we wish to estimate how the region evolves during the collapse and star formation phase. To do so, we adopt the approach of Girichidis et al. (2014) to estimate the free-fall collapse and resultant star formation rate. The mathematics of the approach are described in Appendix B, but we briefly and qualitatively describe the model here.
We start with a gravitationally unstable GMC of radius $R_{\mathrm{SFR}}$, which is initially the largest scale on which a given turbulent trajectory is unstable (Appendix A.2). This region collapses in free-fall. Physically, it does not do so monolithically, but because smaller sub-regions of the cloud may be higher density (shorter free-fall time-scales), they fragment first and continue to accrete from their parent structures (Hoyle, 1953). By assuming that each part of the cloud is in free-fall, we can estimate the fraction of the cloud has collapsed beyond some density threshold to form stars from the initial probability density function for density on the scale $R_{\mathrm{SFR}}$. We define this density threshold by a free-fall time of $0.1$ Myr, although our results are not sensitive to this choice providing it is much shorter than the global collapse time-scale ($\sim 1{-}3$ Myr for our star forming regions – see Figure 11). We emphasise that this does require knowledge of a geometrical density structure during the cloud collapse. We need only assume a geometry when when we compute the evolution of the local density and velocity for individual stars (Section 2.4). We assume the collapse proceeds until the cloud is instantly dispersed when a star formation efficiency $\epsilon=0.5$ is reached, this value being commensurate with simulations of stellar core formation (Matzner & McKee, 2000).
2.3.2 Caveats for the cloud collapse model
@@
Our philosophy in this work is to maintain a model that is as simple as possible, building a framework that can be developed in subsequent work. As a result, we have neglected numerous processes that may influence star formation. In particular, the cloud is exposed to heating by feedback processes from the stars as they form. A wide range of analytic and semi-analytic approaches have been developed to describe this process (e.g. Vázquez-Semadeni et al., 2019), and these may be adopted in future work. We are helped in this context in our interest in low mass star forming regions that are locally well sampled. We do not expect many massive stars or supernovae among these local regions, and we are therefore somewhat justified in neglecting feedback processes and exclusively considering gravitational collapse.
@@ 2.4.1 Approach summary
Having defined a local star formation rate within an individual star forming region (which may form many stars), we can draw formation times for an initial stellar population. Since these stars are initially bound to the cloud from which they form, we integrate the orbit of the star within the gravitational potential of the cloud, ignoring interactions with other stars. For numerical convenience we will calculate the gravitational potential with a smooth Plummer profile with a core radius $a=R_{\mathrm{SFR}}/2.6$, such that the outer radius of the star forming region is twice the half-mass radius. The radius $R_{\mathrm{SFR}}$ evolves following free-fall collapse, as described in Section 2.3 – i.e. $R_{\mathrm{SFR}}$ (and therefore $a$) evolves from its initial value following equation B4, with $r_{1}$ being the initial size of the region. We place stars within the core radius, and we draw an isotropic velocity distribution with a dispersion assuming instantaneous virial equilibrium with the cloud. To track local density and relative velocity of the gas, we then assume the same smooth Plummer density profile (as for Throop & Bally, 2008, see discussion in Section 2.4.2). We then evolve the star under the time-evolving potential, until the final dispersal of the cloud (when $M_{*}=\epsilon M_{\mathrm{tot}}$, for $\epsilon=0.5$). We do not modify the total mass in the region as gas is converted into stars, although given that we end the collapse phase at star formation efficiency $\epsilon=0.5$, this should only have an order unity influence on the total potential.
2.4.2 Caveats for accretion during the bound phase
During the bound phase, apart from our inclusion in the broader star forming context, this part of our approach makes similar simplifications as those of Throop & Bally (2008). As did Throop & Bally (2008), we neglect turbulent fluctuations of the medium within the collapsing region. Clearly gravitational collapse of the ISM modifies the local turbulent energy spectrum, and therefore a prescription for these fluctuations would not be straightforward. We expect our approach to recover typical accretion rates within a given region, but underestimates the variance in these rates. Unlike Throop & Bally (2008), we consider the increase in overall density resulting from the collapse of the star forming region, an appropriate age dispersion of the stars, and a self-consistent time-scale in which stars remain in the bound region.
Another related simplification is that we also do not capture the protostellar/class 0 collapse phase, where for the first $\sim 10^{5}$ years the young system is clearly being fed by the environment (Pineda et al., 2023). We simply form our star instantaneously at some time. In principle, a semi-analytic model that follows gravitational collapse and the influence on the local density distribution is possible (e.g. Vázquez-Semadeni et al., 2024). Such a treatment of the formation stage may be able to self-consistently capture both the initial mass function and produce a physically motivated initial disc from the residual angular momentum. This would again complicate our model, and we therefore do not present a full prescription for these early stages in this work. However, we explore the consequences of our assumed initial conditions for our disc evolution model in Section 3.4.1.
Finally, we do not include the influence of encounters with other stars. This may be included in the future, although for the local star formation regions that are the topic of this work stellar densities are low.
2.5 Ambient density and velocity in the unbound ISM
2.5.1 Approach summary
Once the star forming region is dispersed, all the stars it hosts are instantaneously ejected into the background ISM. The background here is defined by the trajectory which evolves from the initial conditions determined at the moment of collapse of the star forming region. From this moment of collapse, we must then track the density and velocity perturbations at each scale in the trajectory following the procedure of Hopkins (2012), as described in Appendix A.3. At each time-step we (independently) perturb the velocity and logarithmic density perturbation at each scale stochastically, with a magnitude that depends on the turbulent time-scale. This density structure is not used to compute the ambient environment of the star until the collapsing stat forming region is dispersed.
@@ 2.5.1 Approach summary
We continue to evolve the ISM after the dispersal of the cloud and the star moves through the unbound ISM. The relative velocity of the star as it moves through the unbound ISM is the gas velocity at the relevant scale, minus the velocity of the star with respect to the bound cloud at the time of dispersal and the velocity of the cloud itself. The velocity of the cloud is fixed as the gas velocity at the time and scale of cloud collapse. Any star therefore has an associated ambient density and relative velocity at each scale.
2.5.2 Caveats for the turbulent fluctuations of the unbound ISM
The main simplifying assumption for this step is that local star formation does not influence the turbulent fluctuations of the ISM. To varying degrees, this is probably not the case. For example, a local increase in the gravitational potential should drive density enhancements to smaller scales, while stellar feedback drives density enhancements to larger scales. We discuss the role of feedback on the ultimate BHL accretion rates in Sections 3.4.2 and 3.4.3. We expect that our focus on lower mass regions at least somewhat mitigates the exclusion of these influences. Incorporating a prescription for how local star formation shapes the turbulent energy spectrum remains a goal for future work.
@@ 2.6 Accretion rate
2.6.1 Accretion radius
BHL accretion has been the topic of many works (for a review, see Edgar, 2004). In the purest case, it is the gravitational focusing of ambient gas as the star passes through, resulting in a stagnation point in the wake of the star within which material is accreted. The instantaneous BHL radius within which material is captured is then defined:
@@ 2.6.1 Accretion radius
Streamlines with an impact parameter at the BHL radius $R_{\mathrm{BHL}}$ hit a stagnation point along the axis coincident with the relative upstream gas velocity vector. In the ideal case, in an homogeneous medium, they then accrete directly onto the star. In general, material within the cross-section carved out by the accretion radius $R_{\mathrm{acc}}$ (where typically $R_{\mathrm{acc}}=R_{\mathrm{BHL}}$) is captured by the star. At any turbulent scale, we can estimate the instantaneous rate at which a star captures material in such a scenario (Bondi, 1952; Shima et al., 1985):
@@
In some circumstances the radius within which capture can occur can be smaller than the BHL radius, $R_{\mathrm{acc}}<R_{\mathrm{BHL}}$. Broadly, some considerations that may alter the radius from which a star can accrete or suppress accretion entirely are as follows:
  1. 1.
    Tidal radius: The star is not isolated, but embedded in a larger scale gravitational potential. At the tidal radius $R_{\mathrm{T}}$, the tidal force of this larger scale structure balances the force induced by the stellar potential. If the tidal radius $R_{\mathrm{T}}<R_{\mathrm{BHL}}$, then this radius limits the material that can be captured by the star. If a star is bound to a mass $M_{\mathrm{b}}$ at a radius $r_{\mathrm{b}}$, then:
    Initially, this radius is determined by the collapsing cloud, where $r_{\mathrm{b}}$ is the instantaneous location of the star with respect to the centre of the region. Once the cloud has dispersed, we take the galactic values $r_{\mathrm{b}}=8$ kpc and $M_{\mathrm{b}}=10^{12}\,M_{\odot}$.
  2. 2.
    Turbulent scale: Clearly the accretion radius cannot exceed the turbulent size scale under consideration, $R_{\mathrm{t}}$.
  3. 3.
    Photoevaporative wind: High energy photons from the central star can heat up the surface layers of the infalling column and drive a thermal wind, similar to the photoevaporative winds that are driven from the protoplanetary disc surface (e.g. Ercolano & Pascucci, 2017).
  4. 4.
    External irradiation: Stellar feedback from neighbouring stars on scales much larger than the disc may heat the ambient gas, increasing the sound speed and reducing the local density.
In our fiducial model, we do not include photoevaporative suppression or large scale feedback, but discuss them in greater detail in Sections 3.4.2 and 3.4.3. Based on the above considerations, we define the accretion radius:
2.6.2 Accretion histories
In order to compute accretion histories, we draw a representative sample of stars from the representative sample of star forming regions. This involves drawing a star from a region with a probability that is proportional to both the total stellar mass of the star forming region, and at a time that is proportional to the instantaneous star formation rate within each region. It is also now necessary to define the mass of a star to calculate accretion rates. Hereafter we draw stellar masses from a log-uniform distribution between $0.01-3\,M_{\odot}$. This distribution is not important because we will resample when comparing to observations, except that we need to fully sample the stellar mass range relevant for drawing comparisons to protoplanetary disc surveys. Although the stellar masses do not evolve self-consistently as the stars accrete in our model, the total mass that they would gain is nearly always a small fraction of the initial stellar mass. We then draw 500 accretion histories from a representative sample of synthetic star forming regions, with a draw probability weighted by the stellar mass of the region.
@@ 2.6.2 Accretion histories
The resultant BHL accretion rates, normalised for the a solar mass star, are shown in Figure 2, for a subset of six stars in our sample. The phase in which the star remains within the bound gaseous reservoir typically persists for $\sim 1$ Myr. During this phase, local low velocity and high density gas results in an early phase of enhanced BHL accretion at a median rate $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-8}\,M_{\odot}$ yr-1 (Figure 2, bottom). This median rate then quickly drops to $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-10}{-}10^{-11}\,M_{\odot}$ yr-1 in the age range $1{-}3$ Myr for a solar mass star. This is considerably lower than the typical rate of accretion onto the central star, $\dot{M}_{\mathrm{acc}}\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 (Manara et al., 2023). One might therefore assume that BHL accretion no longer plays an important role for disc evolution. However, in a turbulent medium, vorticity of the in-falling material means that accretion is mediated via a (protoplanetary) disc (Krumholz et al., 2006; Kuffmeier et al., 2020). In this case, the change in accretion rate can be averaged over the time-scale $\tau_{\mathrm{acc}}$ at which angular momentum is transported within the disc. Empirically, this time-scale can be a considerable fraction of the disc lifetime (Almendros-Abad et al., 2024). In this case the median underestimates the true contribution of BHL accretion rate onto the disc to the accretion rate onto the star. By contrast, the mean accretion rate when averaged over all stars in our model is $\langle\dot{M}_{\mathrm{BHL}}\rangle\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 at $1-3$ Myr, as shown as the dashed red line in Figure 2. However, the mean overestimates the typical BHL accretion rate contribution for individual discs, since it is dominated by a small number of rapid accretors. Interpreting the contribution from BHL accretion to the stellar accretion rate for typical star-disc systems is therefore not a simple task. Compounding this problem further, discs that survive to comprise the observed sample may also preferentially be those which have been replenished by the ISM. Interpreting the role of BHL accretion for disc evolution therefore requires further calculation, which is the subject of Section 3.
@@ 3.1 Simple disc model
To evolve the disc mass, in our fiducial model we fix the initial $M_{\mathrm{disc}}(t=0)=0$. This is equivalent to the assumption that the initial disc that forms due to core collapse is low mass, or rapidly depleted. This runs counter to the conventional assumption that the protoplanetary disc represents an evolved state of the primordial protostellar disc, which formed during the stellar core collapse. We make this decision as a ‘clean’ test of our model; we wish to explore the properties of discs arising from BHL accretion alone. Such a model requires no additional free parameters. We emphasise that no specific observational constraints contradict a low initial disc mass. Any such evidence would need to connect the primordial circumstellar material, which forms over $\sim 10^{2}$ years, to the protoplanetary disc, which in the context of this work forms over $\sim 10^{5}$ years and evolves over $\sim 10^{6}$ years. Although (massive) discs around Class 0 stars have been inferred (Maury et al., 2019; Tobin et al., 2020) interpreting these observations in terms of the exact age of the (proto-)star (down to $\sim 10^{4}$ years), the nature of the disc and the time-scale over which any such a disc should survive is not simple. Since few, if any, unambiguous observations exist constraining the properties of primordial discs that are categorically around stars of age $\ll 10^{5}$ years, there is no reason to assume that such a primordial disc should be both long-lived and comparable in mass to the subsequently accreted one. Indeed, it would be a remarkable coincidence if BHL accretion and protostar formation contribute an equal fraction of mass to the disc population. Nonetheless, we also discuss the impact of this assumption on our results in Section 3.4.1.
@@
For $\tau_{\mathrm{acc}}=M_{\mathrm{disc}}/\dot{M}_{\mathrm{acc}}$ we assume a log-uniform distribution ±$\pm 1$ dex around $\log\tau_{\mathrm{acc}}/\mathrm{1\,Myr}=-0.5$, based on visual inspection of the observed distribution (Almendros-Abad et al., 2024). Empirically, the time-scale $\tau_{\mathrm{acc}}$ is not strongly correlated with stellar mass, but may be a complex function of the disc properties or stellar age. For example, visually interpreting Figure 12 of Almendros-Abad et al. (2024), the median accretion time-scale appears to be $\tau_{\mathrm{acc,1/2}}\sim 0.1$ Myr for the young region $\rho$ Ophiuchus, while $\tau_{\mathrm{acc,1/2}}\sim 1$ Myr for the older Lupus region. Indeed, we expect rapid BHL accretion to instigate turbulence in the disc, temporarily decreasing $\tau_{\mathrm{acc}}$. However, empirically it is not clear the degree to which any variation of $\tau_{\mathrm{acc}}$ is a survival bias, discs with shorter $\tau_{\mathrm{acc}}$ having been depleted (we will show a similar evolution in our model which has fixed $\tau_{\mathrm{acc}}$ for each star in Section 3.2.2). While exploring time- or in-fall-dependent $\tau_{\mathrm{acc}}$ is an interesting avenue for future exploration, we initially maintain our simplified model, subsequently exploring the potential role of the ISM in driving a time-dependent stellar accretion rate (Section 4.1).
@@
Given the initial condition and accretion time-scale, we evolve each disc mass according to equation 8 over the entire time period under consideration. However, we assume a simple criterion for disc dispersal: $M_{\mathrm{disc}}<3\times 10^{-5}\,M_{\odot}$. We adopt this criterion when showing disc properties in subsequent analysis. This threshold is motivated empirically by the fact that few discs have masses below this value; our results are not strongly sensitive to this choice within a sensible range (see Appendix C).